a square is changed into a rectangle by increasing its length by 10%.how is the area changed?

Found 2 solutions by sudhanshu_kmr, Alan3354:
Answer by sudhanshu_kmr(1152)

Let side of square = x
area = x^2
new area = x^2* 110/100

so, area increased by 10%.

Answer by Alan3354(69443)

If only one dimension is changed, the area is increased by 10%.
---------
Area = w*L original
Area = w*1.1L = 1.1*original area